OpenStudy (anonymous):

How many different positive 3-digit integers can be formed if the three digits four, five, and six must be used in each of the integers? 3 4 6 8 9

OpenStudy (anonymous):

first place can be filled in 3 ways (4,5,6) another one is same(4,5,6) and third one is also (4,5,6) so totally as they are dependent so total no. ways are 3*3*3=27 but here we have to just 3 nos and three places all have to come so only 3*2(first place 3 second 2)=6 ways is the answer (456,546,654,465,564,645)....:)

OpenStudy (anonymous):

Can the Digits Be repeated....if Yes then it is 3^3 = 27 if not then it is 3! = 3*2 * 1 =6